Simplify: (1 + i)/(2 - i).
Thanks

Found 2 solutions by fractalier, Nate:
Answer by fractalier(6550)

We need to multiply top and bottom by the conjugate of the bottom...here that's 2+i, so we get
(1 + i)/(2 - i) =
(1 + i)(2 + i) / (2 - i)(2 + i) =
(2 + i + 2i - 1) / (4 + 2i - 2i + 1) =
(1 + 3i) / 5

Answer by Nate(3500)